# Multiple-policy Evaluation via Density Estimation

### v1

@@ A.2 Proof of the optimal sampling distribution (4)
$|\hat{V}^{\pi^{k}}_{1}-V^{\pi^{k}}_{1}|\leq\sqrt{\frac{2H\rho_{\mu,k}\log(1/\delta)}{n}}+\frac{2M_{k}\log(1/\delta)}{3n}$
@@
$n_{\mu,k}\leq\frac{8H\rho_{\mu,k}\log(1/\delta)}{\epsilon^{2}}+\frac{4M_{k}\log(1/\delta)}{3\epsilon}$
@@
$n_{\mu}\leq\frac{8H\max_{k\in[K]}\rho_{\mu,k}\log(K/\delta)}{\epsilon^{2}}+\frac{4M\log(K/\delta)}{3\epsilon}$
@@ A.3 An example of unrealizable optimal sampling distribution
The gradient of $\ell^{\pi}_{h}(w)$,
$\nabla_{w(s,a)}\ell^{\pi}_{h}(w)=\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}w(s,a)-\sum_{s^{\prime},a^{\prime}}\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})P(s|s^{\prime},a^{\prime})\pi(a|s)\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}$
@@
$=$ $\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi}_{h}(s,a)\right|-\left|\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\pi(a|s)\left(d^{\pi}_{h-1}(s^{\prime},a^{\prime})-\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}\right)\right|$
@@
$\sum_{s,a}\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi}_{h}(s,a)\right|$ $\leq\epsilon+\sum_{s,a}\left|\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\pi(a|s)\left(d^{\pi}_{h-1}(s^{\prime},a^{\prime})-\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}\right)\right|$
@@
$w^{*}_{h}(s,a)=\frac{d^{\pi}_{h}(s,a)}{\tilde{\mu}_{h}(s,a)}\hat{\mu}_{h}(s,a)\leq\frac{\frac{4}{3}\hat{d}^{\pi}_{h}(s,a)}{\frac{4}{5}\hat{\mu}_{h}(s,a)}\hat{\mu}_{h}(s,a)=\frac{5}{3}\hat{d}^{\pi}_{h}(s,a)$
@@
$g_{h}(w)(s,a)=\frac{w(s,a)}{\hat{\mu}_{h}(s,a)}\mathbb{I}(s^{i}_{h}=s,a^{i}_{h}=a)-\frac{\hat{w}_{h-1}(s^{j}_{h-1},a^{j}_{h-1})}{\hat{\mu}_{h-1}(s^{j}_{h-1},a^{j}_{h-1})}\pi(a|s)\mathbb{I}(s^{j}_{h}=s)$
the variance bound is,
@@
$\leq O\left(\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}+\frac{(\hat{d}_{h-1}^{\pi}(s,a))^{2}}{\hat{\mu}_{h-1}(s,a)}\right)$ (16)
the last inequality is due to the bounded feasible set for $w$ and the property of coarse distribution estimator $\tilde{\mu}_{h}(s,a)\leq\frac{4}{3}\hat{\mu}_{h}(s,a)$.
@@
$\sum_{s,a}\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi}_{h}(s,a)\right|\leq\frac{\epsilon}{4H},\forall h=1,2,\dots,H$
@@
Invoke the convergence rate for strongly-convex and smooth loss functions, i.e. Lemma 4.6, we have that the number of samples needed to achieve $\ell^{\pi}_{h}(\hat{w}_{h})-\ell^{\pi}_{h}(w^{*}_{h})\leq\frac{\epsilon^{2}}{32\xi H^{4}}$ is,
$n=O\left(\frac{\xi}{\gamma}\frac{H^{4}G^{2}}{\epsilon^{2}}\right)$
We have shown in (8) that $\frac{\xi}{\gamma}\leq\frac{5}{3}$, this nice property helps us to get rid of the undesired ratio of the smoothness factor and the strongly-convexity factor, i.e. $\frac{\max_{s,a}\mu(s,a)}{\min_{s,a}\mu(s,a)}$ of the original loss function (7) which can be extremely bad. Replacing $G^{2}$ by our variance bound (A.5), we have,
$n^{\pi}_{h}=O\left(\frac{H^{4}}{\epsilon^{2}}\left(\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}+\frac{(\hat{d}_{h-1}^{\pi}(s,a))^{2}}{\hat{\mu}_{h-1}(s,a)}\right)\right)$
@@
$n^{\pi}=O\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}\right)$
@@
$\mathbb{P}(|\hat{\mu}_{MoM}-\mu|>\epsilon)$ $\leq\mathbb{P}(\sum_{i=1}^{N}\mathbb{I}(|\hat{\mu}_{i}-\mu|>\epsilon)\geq\frac{N}{2})$

### v2

@@ A.2 Derivation of the optimal sampling distribution (4)
$|\hat{V}^{\pi^{k}}_{1}-V^{\pi^{k}}_{1}|\leq\sqrt{\frac{2H\rho_{\mu,k}\log(1/\delta)}{n}}+\frac{2M_{k}\log(1/\delta)}{3n},$
@@
$n_{\mu,k}\leq\frac{8H\rho_{\mu,k}\log(1/\delta)}{\epsilon^{2}}+\frac{4M_{k}\log(1/\delta)}{3\epsilon}.$
@@
$n_{\mu}\leq\frac{8H\max_{k\in[K]}\rho_{\mu,k}\log(K/\delta)}{\epsilon^{2}}+\frac{4M\log(K/\delta)}{3\epsilon},$
@@ A.3 An example of unrealizable optimal sampling distribution
The gradient of $\ell^{\pi}_{h}(w)$ is,
$\nabla_{w(s,a)}\ell^{\pi}_{h}(w)=\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}w(s,a)-\sum_{s^{\prime},a^{\prime}}\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})P(s|s^{\prime},a^{\prime})\pi(a|s)\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}.$
@@
$=$ $\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi}_{h}(s,a)\right|$
$-\left|\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\pi(a|s)\left(d^{\pi}_{h-1}(s^{\prime},a^{\prime})-\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}\right)\right|.$
@@
$\sum_{s,a}$ $\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi}_{h}(s,a)\right|$
$\leq\epsilon+\sum_{s,a}\left|\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\pi(a|s)\left(d^{\pi}_{h-1}(s^{\prime},a^{\prime})-\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}\right)\right|$
@@
$w^{*}_{h}(s,a)=\frac{d^{\pi}_{h}(s,a)}{\tilde{\mu}_{h}(s,a)}\hat{\mu}_{h}(s,a)\leq\frac{\frac{4}{3}\hat{d}^{\pi}_{h}(s,a)}{\frac{4}{5}\hat{\mu}_{h}(s,a)}\hat{\mu}_{h}(s,a)=\frac{5}{3}\hat{d}^{\pi}_{h}(s,a).$
@@
$g_{h}(w)(s,a)=\frac{w(s,a)}{\hat{\mu}_{h}(s,a)}\mathbb{I}(s^{i}_{h}=s,a^{i}_{h}=a)-\frac{\hat{w}_{h-1}(s^{j}_{h-1},a^{j}_{h-1})}{\hat{\mu}_{h-1}(s^{j}_{h-1},a^{j}_{h-1})}\pi(a|s)\mathbb{I}(s^{j}_{h}=s).$
The variance bound becomes
@@
$\leq O\left(\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}+\frac{(\hat{d}_{h-1}^{\pi}(s,a))^{2}}{\hat{\mu}_{h-1}(s,a)}\right),$ (15)
where the last inequality is due to the bounded feasible set for $w$ and the property of coarse distribution estimator $\tilde{\mu}_{h}(s,a)\leq\frac{4}{3}\hat{\mu}_{h}(s,a)$.
@@
$\sum_{s,a}\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi}_{h}(s,a)\right|\leq\frac{\epsilon}{4H},\forall h=1,2,\dots,H,$
@@
Lemma A.4.
$\mathbb{E}[L(\hat{w})-L(w^{*})]\leq\frac{2G^{2}}{\lambda(T+1)},$
where $G^{2}$ is the variance bound of the stochastic gradient.
Invoke the convergence rate for strongly-convex and smooth loss functions, i.e. Lemma A.4, we have that the number of samples needed to achieve $\ell^{\pi}_{h}(\hat{w}_{h})-\ell^{\pi}_{h}(w^{*}_{h})\leq\frac{\epsilon^{2}}{32\xi H^{4}}$ is,
$n=O\left(\frac{\xi}{\gamma}\frac{H^{4}G^{2}}{\epsilon^{2}}\right).$
We have shown in Section 4.3 that $\frac{\xi}{\gamma}\leq\frac{5}{3}$, this nice property helps us to get rid of the undesired ratio of the smoothness factor and the strongly-convexity factor, i.e. $\frac{\max_{s,a}\mu(s,a)}{\min_{s,a}\mu(s,a)}$ of the original loss function (7) which can be extremely bad. Replacing $G^{2}$ by our variance bound (A.5), we have,
$n^{\pi}_{h}=O\left(\frac{H^{4}}{\epsilon^{2}}\left(\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}+\frac{(\hat{d}_{h-1}^{\pi}(s,a))^{2}}{\hat{\mu}_{h-1}(s,a)}\right)\right).$
@@
$n^{\pi}=O\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}\right).$
@@
$\mathbb{P}(|\hat{\mu}_{MoM}-\mu|>\epsilon)$ $\leq\mathbb{P}(\sum_{i=1}^{N}\mathbb{I}(|\hat{\mu}_{i}-\mu|>\epsilon)\geq\frac{N}{2}).$